OpenStudy (experimentx):

Evaluate: \[ \displaystyle \int_0^{\infty} \dfrac{(\log x)^2}{x^2 + 1} dx \]

OpenStudy (experimentx):

http://tinyurl.com/9csdea3

OpenStudy (anonymous):

*

OpenStudy (anonymous):

going with \[t=\log x\]\[\displaystyle \int_{-\infty}^{\infty} \dfrac{t^2e^t}{1+ e^{2t}} \text{d}t=\int_{-\infty}^{0} \dfrac{t^2e^t}{1+ e^{2t}} \text{d}t+\int_{0}^{\infty} \dfrac{t^2e^t}{1+ e^{2t}} \text{d}t\]first one for example\[\int_{-\infty}^{0} \dfrac{t^2e^t}{1+ e^{2t}} \text{d}t=\int_{-\infty}^{0} t^2e^t \sum_{n=0}^{\infty}(-1)^ne^{2nt} \text{d}t= \sum_{n=0}^{\infty} (-1)^n\int_{-\infty}^{0} t^2e^te^{2nt} \text{d}t\\=\sum_{n=0}^{\infty} (-1)^n \frac{2}{(2n+1)^3}\]

OpenStudy (anonymous):

but i'd like to know how can we go with complex integration

OpenStudy (experimentx):

the series looks like Fourier expansion.

OpenStudy (anonymous):

yeah

OpenStudy (experimentx):

hold on .. is that 3?

OpenStudy (anonymous):

lol .. yes

OpenStudy (anonymous):

this one is reachable by fourier series i think

OpenStudy (experimentx):

you made things a bit more complicated http://www.wolframalpha.com/input/?i=sum [1%2F%282n%2B1%29^3%2C+{n%2C+0%2C+Infinity}]

OpenStudy (experimentx):

but this seems nice http://www.wolframalpha.com/input/?i=sum [%28-1%29^n%2F%282n%2B1%29^3%2C+{n%2C+0%2C+Infinity}]

OpenStudy (anonymous):

i made it worser

OpenStudy (experimentx):

No not really ,,, this is nice ... and interesting http://www.wolframalpha.com/input/?i=sum [%28-1%29^n%2F%282n%2B1%29^3%2C+{n%2C+1%2C+Infnity}]

OpenStudy (anonymous):

ahh yeah so this is reachable man

OpenStudy (experimentx):

probably some nasty Fourier analysis.

OpenStudy (anonymous):

*

OpenStudy (anonymous):

Hae you tried using the residue theorem?

OpenStudy (experimentx):

tried ... but stuck. I need picture of contour .... that t^2 term is bugging me badly http://openstudy.com/users/experimentx#/updates/50476f64e4b0c3bb09860ba6

OpenStudy (anonymous):

See http://en.wikipedia.org/wiki/Methods_of_contour_integration Example (V)

OpenStudy (experimentx):

thanks ... i this is helpful ...

OpenStudy (experimentx):

*think

OpenStudy (experimentx):

i'll try it ... if it get answer i'll post solution.

OpenStudy (anonymous):



OpenStudy (anonymous):

man this contour works

OpenStudy (experimentx):

did you try it?

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

oh great ... but I have QM exam six days later.

OpenStudy (anonymous):

\[\oint \frac{z^2 e^z}{1+e^{2z}} dz=a_{-1}\]

OpenStudy (anonymous):

man try it after exams

OpenStudy (experimentx):

sure ... 14 days to go.